lemon1boy1:

The equation of line LM is 5x − y = −4. What is the equation of a line perpendicular to line LM in slope-intercept form that contains point (−3, 2)?

aevans182:

2=-1/5x-3

lemon1boy1:

The answer choices are ' y = 5x + 13 y = negative one fifth x + seven fifths y = negative one fifth x − seven fifths y = 5x − 17

rocfp:

The equation of line LM is 5x − y = −4. What is the equation of a line perpendicular to line LM in slope-intercept form that contains point (−3, 2)? The equation of line LM is 5x − y = −4. What is the equation of a line perpendicular to line LM in slope-intercept form that contains point (−3, 2)? Since the product of slopes of two perpendicular lines is -1. Hence, the equation of a line perpendicular to line LM in slope-intercept form that contains the point (−3, 2) is . Hence, the correct answer is b) y = negative one fifthx + seven fifths I hope this help :3

lemon1boy1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rocfp The equation of line LM is 5x − y = −4. What is the equation of a line perpendicular to line LM in slope-intercept form that contains point (−3, 2)? The equation of line LM is 5x − y = −4. What is the equation of a line perpendicular to line LM in slope-intercept form that contains point (−3, 2)? Since the product of slopes of two perpendicular lines is -1. Hence, the equation of a line perpendicular to line LM in slope-intercept form that contains the point (−3, 2) is . Hence, the correct answer is b) y = negative one fifthx + seven fifths I hope this help :3 \(\color{#0cbb34}{\text{End of Quote}}\) thanks

rocfp:

NP :)